OpenStudy (anonymous):

determine the value: 10% of $365

OpenStudy (anonymous):

"%" means "/100" (as in per each hundred) "of" means "times" (as in multiplication) so: 10% of $365 is (10/100) * $365 =(1/10) * $365 =$365/10 =$36.5

OpenStudy (anonymous):

thank-you !

OpenStudy (anonymous):

you are most welcome.. it helps to remember this: "%" means "/100" (as in per each hundred) "of" means "times" (as in multiplication)